OpenStudy (anonymous):

what is the ordered pair? -2x-3y=19 -5,-4 -2,-5 -3,-5 -5,-2??? idk what a ordered pair is...

OpenStudy (anonymous):

@MinhLe pleasee helppppp!

OpenStudy (anonymous):

We have the function : -2x-3y Then you put these number to x and y and calculate . Example x=-5 , y=-4 : We have -2(-5) -3(-4) = 10+12 = 22 # 19 => this isn't the ordered pair . Similar to the orders .

OpenStudy (anonymous):

i kinda need an answerr..............im checking my friends answer

OpenStudy (anonymous):

@MinhLe